OpenStudy (anonymous):

Complete the equation for the inverse of f(x) F(x)= (3x-2)/6 F^-1(y)=????

OpenStudy (aravindg):

Put f(x)=y Then solve for x.

OpenStudy (anonymous):

I don't understand...

OpenStudy (anonymous):

Switch x and y and solve for y.

OpenStudy (anonymous):

f(x) = (3x-2)/6 f(x) = y y = (3x-2)/6 now switch x and y and solve for y to find the inverse ^^

OpenStudy (anonymous):

I feel so stupid for not understanding..... It's just going against what the lesson told me.... I think that's all.

OpenStudy (anonymous):

The lesson basically says to just switch everything .... /:

OpenStudy (aravindg):

It's okay..You can learn something new here :)

OpenStudy (aravindg):

y = (3x-2)/6 x=?

OpenStudy (anonymous):

I have no idea... Because how would I know if I didn't know what y was?

OpenStudy (anonymous):

f(x) is another way to write y, it's just a different notation ^^

OpenStudy (anonymous):

Okay.... So how would I know what x is? Ugh I'm so confused /: I'm sorry to both of you. This is all new to me.

OpenStudy (aravindg):

Okay I have a better method if u find this confusing.

OpenStudy (anonymous):

Wait it's asking for f^-1(y)???

OpenStudy (anonymous):

Yes it is.

OpenStudy (aravindg):

y = (3x-2)/6 x=?

OpenStudy (anonymous):

I'm trying to find f^-1(y) the inverse to f(x)...

OpenStudy (aravindg):

yes so answer my qn. x=?

OpenStudy (anonymous):

It doesn't tell me...

OpenStudy (aravindg):

Find x!!!

OpenStudy (anonymous):

Hahaha, re-arrange the equation you have for x--so try to rearrange and leave x on one side using your algebra skills ^^

OpenStudy (anonymous):

I don't know how to do this.... I feel so freaking stupid lol. It's because of the division /: